# CodeWiki: Evaluating AI’s Ability to Generate Holistic Documentation for Large-Scale Codebases

### v3

@@ 1 Introduction
$\hookrightarrow$ To address these challenges, we present CodeWiki, the first open-source framework specifically designed to generate holistic, structured repository-level documentation that includes diverse content types such as system architecture diagrams, data flow visualizations, and sequence diagrams while scaling effectively for projects of arbitrary size. CodeWiki introduces several key innovations: a hierarchical decomposition strategy inspired by dynamic programming principles that breaks complex repositories into manageable modules while preserving architectural coherence, a recursive agentic system with dynamic delegation capabilities that enables adaptive processing based on module complexity, and a comprehensive documentation synthesis approach that generates multiple forms of visual and textual documentation reflecting fine-grained codebase collaboration patterns.
@@
$\hookrightarrow$ To fill this gap, we introduce CodeWikiBench, the first benchmark specifically designed for repository-level documentation. Instead of relying on surface-level similarity, we derive hierarchical rubrics from the official documentation maintained by project developers, enabling the evaluation of underlying repository properties. For robustness, CodeWikiBench further adopts an agentic assessment framework in which judge agents evaluate generated documentation against rubric requirements and aggregate results hierarchically, producing both quality scores and reliability measures.
@@
    CodeWiki, the first open-source semi-agentic framework for scalable repository-level documentation with proprietary and open-source LLM backends.
@@
    CodeWikiBench, the first repository-level documentation benchmark with hierarchical rubrics and agentic assessment.
@@
    First multilingual repository documentation supporting 7 widely used languages.
@@ 3.3 Hierarchical Assembly and Documentation Synthesis
4 The First Benchmarking System - CodeWikiBench
@@ 5.2 Baseline
To the best of our knowledge, CodeWiki represents the first open-source framework specifically designed for comprehensive repository-level documentation generation. The absence of comparable open-source solutions necessitates comparison with closed-source alternatives. We select DeepWiki [10] as our primary baseline, as it shares our goal of automated repository-level documentation generation and has demonstrated effectiveness in industrial applications.
@@ 5.4.1 RQ1: Documentation Quality and Coverage
Answer to RQ1: CodeWiki demonstrates superior overall documentation quality and coverage compared to existing systems. With proprietary models, it achieves 4.73% improvement; with open-source models, it matches baseline performance while offering transparency and accessibility benefits.
@@ 5.4.2 RQ2: Cross-Language Generalization
CodeWiki-sonnet-4 demonstrates exceptional performance on high-level scripting languages, achieving 79.14% average across Python, JavaScript, and TypeScript repositories—a 10.47 percentage point improvement over DeepWiki’s 68.67%. CodeWiki-kimi-k2 achieves 77.31% on these languages, showing strong performance with open-source models (+8.64% over DeepWiki).
@@
Answer to RQ2: CodeWiki demonstrates strong cross-language generalization with exceptional performance on high-level languages (79.14% with proprietary models, 77.31% with open-source models) and consistent improvements on managed languages, while systems programming languages present similar challenges across all frameworks and model types.
@@ Appendix D Complementary Experimental Results
These complementary results reinforce our findings from the main experiments, showing an average improvement of 1.25% across 14 diverse repositories. Notably, CodeWiki achieves substantial gains in several large-scale projects, including a remarkable 29.53% improvement on the Trino database engine (1.5M LOC) and 11.42% on the Storybook UI development platform (365K LOC). The results further validate that CodeWiki’s hierarchical approach generalizes well across different repository characteristics and programming paradigms.

### v4

@@ 1 Introduction
$\hookrightarrow$ To address these challenges, we present CodeWiki, an open-source framework specifically designed to generate holistic, structured repository-level documentation that includes diverse content types such as system architecture diagrams, data flow visualizations, and sequence diagrams while scaling effectively for projects of arbitrary size. CodeWiki introduces several key innovations: a hierarchical decomposition strategy inspired by dynamic programming principles that breaks complex repositories into manageable modules while preserving architectural coherence, a recursive agentic system with dynamic delegation capabilities that enables adaptive processing based on module complexity, and a comprehensive documentation synthesis approach that generates multiple forms of visual and textual documentation reflecting fine-grained codebase collaboration patterns.
@@
$\hookrightarrow$ To fill this gap, we introduce CodeWikiBench, a comprehensive benchmark specifically designed for repository-level documentation. Instead of relying on surface-level similarity, we derive hierarchical rubrics from the official documentation maintained by project developers, enabling the evaluation of underlying repository properties. For robustness, CodeWikiBench further adopts an agentic assessment framework in which judge agents evaluate generated documentation against rubric requirements and aggregate results hierarchically, producing both quality scores and reliability measures.
@@
    CodeWiki, an open-source semi-agentic framework for scalable repository-level documentation with proprietary and open-source LLM backends.
@@
    CodeWikiBench, a comprehensive repository-level documentation benchmark with hierarchical rubrics and agentic assessment.
@@
    Comprehensive multilingual repository documentation supporting 7 widely used languages.
@@ 3.3 Hierarchical Assembly and Documentation Synthesis
4 Benchmarking System - CodeWikiBench
@@ 5.2 Baseline
CodeWiki provides an open-source framework specifically designed for comprehensive repository-level documentation generation. Due to the limited availability of comparable open-source solutions, we compare with closed-source alternatives. We select DeepWiki [10] as our primary baseline, as it shares our goal of automated repository-level documentation generation and has demonstrated effectiveness in industrial applications.
@@ 5.4.1 RQ1: Documentation Quality and Coverage
Answer to RQ1: CodeWiki demonstrates strong overall documentation quality and coverage compared to existing systems. With proprietary models, it achieves 4.73% improvement; with open-source models, it matches baseline performance while offering transparency and accessibility benefits.
@@ 5.4.2 RQ2: Cross-Language Generalization
CodeWiki-sonnet-4 demonstrates strong performance on high-level scripting languages, achieving 79.14% average across Python, JavaScript, and TypeScript repositories—a 10.47 percentage point improvement over DeepWiki’s 68.67%. CodeWiki-kimi-k2 achieves 77.31% on these languages, showing strong performance with open-source models (+8.64% over DeepWiki).
@@
Answer to RQ2: CodeWiki demonstrates strong cross-language generalization with high performance on high-level languages (79.14% with proprietary models, 77.31% with open-source models) and consistent improvements on managed languages, while systems programming languages present similar challenges across all frameworks and model types.
@@ Appendix D Complementary Experimental Results
These complementary results reinforce our findings from the main experiments, showing an average improvement of 1.25% across 14 diverse repositories. Notably, CodeWiki achieves substantial gains in several large-scale projects, including a 29.53% improvement on the Trino database engine (1.5M LOC) and 11.42% on the Storybook UI development platform (365K LOC). The results further validate that CodeWiki’s hierarchical approach generalizes well across different repository characteristics and programming paradigms.